OpenStudy (anonymous):

Find the least common multiple of x³ – x² + x – 1 and x² – 1 . Write the answer in factored form. becoming FAN AND MEDAL

OpenStudy (anonymous):

(x+1) and (x-1)

OpenStudy (anonymous):

how would you use synthetuc divion?

OpenStudy (anonymous):

so divide x+1 and x+2 by the cubic?

OpenStudy (yacoub1993):

x^3 - x^2 + x - 1 = (x - 1)(x^2 + 1) That is the big key to the problem and that was done using synthetic division. x^2 - 1 = (x - 1)(x + 1) So, you take (x - 1)(x^2 + 1) from the first, but you only have to take "x + 1" from the second because you already have a "x - 1" factor from the first. And that's it!